whats the equation of the line that passes through the given point and is parallel to the given equation. (7,4); 3x-8y=5
Answer by jim_thompson5910(35256)

3x-8y=C

3(7)-8(4)=C ... plug in the point (7,4)

21 - 32 = C

-11 = C

C = -11

So the answer, in standard form, is 3x - 8y = -11